There are a total 108 offensive players on a football team. There are twice as many runners than passing players. Write two equations and then slove for both equation. How many runners are there and how many passing people are there?
Answer by stanbon(75887)

There are a total 108 offensive players on a football team. There are twice as many runners than passing players. Write two equations and then solve for both equation. How many runners are there and how many passing people are there?
---
Equations:
r + p = 108 players
r = 2p
----
Substitute for "r" and solve for "P":
2p + p = 108
3p = 108
p = 36 (# of passers)
----
solve for "r":
r + p = 108
r + 36 = 108
r = 72 (# of runners)
========================
Cheers,
Stan H.